# Volume and topology of bounded and closed hyperbolic $3$-manifolds, II

### v1

@@ 3.2. Collars
We now turn our attention to embedding collars, and controlling their interaction with muffins. The definition of collar given below is implicitly used in [27] and [14]. To motivate it, we recall from Section 2.1 that for a hyperbolic $3$-manifold $N$ with totally geodesic boundary, we take the universal cover $\widetilde{N}$ of $N$ to be a convex subset of $\mathbb{H}^{3}$ bounded by a collection of geodesic hyperplanes. For a component $\Pi$ of $\partial\widetilde{N}$, let $\pi$ be the orthogonal projection to $\Pi$ from Definitions and Remarks 3.1.2. Then for $x\in\Pi$, $\pi^{-1}(x)$ is a geodesic intersecting $\Pi$ orthogonally at $x$, and $\pi^{-1}(x)\cap\widetilde{N}$ is either a ray with endpoint $x$ or a segment with one endpoint at $x$ and the other on a different component of $\partial\widetilde{N}$.
@@
Let $N$ be a hyperbolic $3$-manifold with totally geodesic boundary. If $x$ is a point of $\partial N$ such that there exists a (necessarily unique) geodesic path in $N$ (parametrized by arclength) which begins at $x$, is perpendicular to $\partial N$ at its initial point, and has terminal point in $\partial N$, we shall denote the length of this path by $h_{x}$, the interval $[0,h_{x}]$ by $J_{x}$, and the path itself by $\alpha_{x}:J_{x}\to N$. If $x$ is a point of $\partial N$ for which no path of this type exists, then there is a unique geodesic ray beginning at $x$; in this case, we shall set $h_{x}=+\infty$ and $J_{x}=[0,\infty)$, and denote (the arclength parametrization of) the ray by $\alpha_{x}:[0,\infty)\to N$. For a subset $S$ of $\partial N$, we define the open collar of $S$ in $N$ with height $h>0$ to be the set $\bigcup_{x\in S}\alpha_{x}([0,\min(h,h_{x})))$. We shall say that the collar is embedded if $h\leq\inf\{h_{x}\,|\,x\in S\}$, and the map from $S\times[0,h]$ to $N$ defined by $(x,t)\mapsto\alpha_{x}(t)$ is one-to-one.
@@
The claim implies that the $h$ open collar of the preimage in $\Pi$ of $\partial N-(C\cup C^{\prime})$ is contained in $\widetilde{N}$ and does not intersect the $h$ open collar of any other component of $\partial\widetilde{N}$. Therefore the universal cover induces an embedding of the $h$ open collar of $\partial N-(C\cup C^{\prime})$ in $N$.
We now recall from Fact 3.1.7 that $U_{1}=A+W$. Here $U_{1}$, defined in Lemma 3.1.5, is a lower bound on the distance between a lift of $\lambda_{1}$ and any component of $\partial\widetilde{N}$ not containing either of its endpoints. Since $\mathrm{Muf}_{\ell_{1}}$ is contained in the $W$-neighborhood of its center, and the collar height $h$ is less than $A$, the $h$ open collar of $\Pi$ in $\widetilde{N}$ thus does not intersect any lift of $M$ that has no cap on $\Pi$. This implies that the $h$ open collar of $\partial N-(C\cup C^{\prime})$ in $N$ is disjoint from $M$. ∎
@@
Recall from the proof of Lemma 3.1.8 that $\widetilde{M}$ is contained in a radius-$W_{2}$ neighborhood of its center $\tilde{\lambda}_{2}$, where $W_{2}$ is its waist radius, and from Lemma 3.1.5 that the quantity $U_{2}$ defined in that lemma is a lower bound for the distance from $\tilde{\lambda}_{2}$ to $\Pi_{3}$. From the proof of Proposition 3.1.11, for $(x,y)$ in the rectangle $[a,b]\times[c,d]$, $W_{2}$ is bounded above by $W^{0}_{S}$ and $U_{2}$ is bounded below by $U_{2}^{0}$. Therefore by the triangle inequality, no point lies in both $\widetilde{M}$ and a height-$H^{0}$ open collar  of $\Pi_{3}$. ∎
@@ 4. Volume bounds
[…]
@@
[…]
@@
[…]
@@
[…]
Since $\cosh R$ decreases with $x$, this quantity is positive as long as $2A<2H+\sinh(2H)$. For $x\leq 1.5$ we have $1.52A<\cosh^{-1}(2.34)<$ 1.5 and for $H\geq\cosh^{-1}(1.5)/2$,
$2.082H+\sinh(2H)\geq\cosh^{-1}(1.5)+\sqrt{1.5^{2}-1}>$ 2.08 .
@@
Table 1. Values of $V(x,H)$, truncated after $3$ decimal places, for $H=\cosh^{-1}(Y(x))/2$.
@@
[…]
@@
[…]
@@
[…]
@@
On each maximal subinterval $[a_{0},b_{0})$ of $[1.24,1.5)$ where $Y$ is constant, we find the minimum value of the function $\max\{\ell_{1},E,M\}$ from Lemma 2.2.2. Let $c_{0}$ be the hyperbolic cosine of this value, truncated after three decimal places. By Lemma 2.2.2, for $N$ with $\cosh\ell_{1}\in[a_{0},b_{0}]$, $\cosh\ell_{2}\geq c_{0}$. Let $d_{0}$ be the value of $Y$ on $[a_{0},b_{0})$. We have written Python scripts that divide the rectangle $[a_{0},b_{0}]\times[c_{0},d_{0}]$ into subrectangles of size $0.001\times 0.001$, and on each subrectangle $[a,b]\times[c,d]$ performs the following tasks.
@@
    Check whether $[a,b]\times[c,d]$ intersects the possible subregion: the subset of $[a_{0},b_{0}]\times[c_{0},d_{0}]$ consisting of $(x,y)$-values that satisfy the criterion of Lemma 2.2.2. For instance, $E(x)=\max\{\ell_{1},E,M\}$ for $x\leq 1.366$, so since $E$ is a decreasing function of $x$ here (cf. [14, Lemma 3.4]), for $b\leq 1.366$ this is tantamount to checking that $d\geq\cosh E(b)$.
    If $[a,b]\times[c,d]$ does not intersect the possible subregion then we ignore it since in this case, by Lemma 2.2.2, there does not exist a manifold satisfying the hypotheses of the present result and with an $(x,y)$-value in $[a,b]\times[c,d]$.
@@
    If $[a,b]\times[c,d]$ does intersect the possible subregion, check the muffin-embedding criteria of Proposition 3.1.11. These hold in each such case, confirming the expectation recorded above that result.
@@
[…]
@@
The resulting volume bounds are summarized in Figure 4.2. Each big rectangle $[a_{0},b_{0}]\times[c_{0},d_{0}]$ from above is shaded in the Figure, and the smallest of the volume bounds produced by the process above—taken over all subrectangles $[a,b]\times[c,d]$—is recorded in or directly above the big rectangle. (The subrectangle on which the minimum is attained is also blackened in the Figure.) Since each of these bounds is larger than $7.4$, the Theorem follows from Proposition 4.2. ∎
@@
We use a single large rectangle $[1.23,1.24]\times[1.74,1.938]$. The lower bound of $y=1.74$ is the minimum of $\cosh E(x)$ on this interval, which is attained at $x=1.24$, truncated after three decimal places. The upper bound is the maximum of $Y_{0}(x)$ on the interval, also attained at $1.24$, rounded up after three decimal places. Because $Y_{0}(x)$ is not constant here but increasing, we alter step (1) above for a subrectangle $[a,b]\times[c,d]$ by checking that both $d\geq\cosh E(b)$ and $c\leq Y_{0}(x)$.
To achieve a slight improvement in our volume bounds, we use a subrectangle size of $0.0005\times 0.001$—that is, we divide the previous subrectangles in half vertically. The smallest bound resulting from Theorem 4.5’s revised process is $7.409$, truncated after three decimal places, attained on the subrectangle $[1.2305,1.231]\times[1.808,1.809]$, proving the result. (Using the previous subrectangle size would yield $7.399$ as a bound.) The relevant Python script is VolScript_EUl2.py, and its output is 123_124.txt. ∎
@@ 5. Trimonic Manifolds from a $(1,1,1)$-hexagon
As in the previous proof, we take $\Pi^{\prime}$ to be the component of $\partial\widetilde{N}$ containing the external edge of $C$ closest to $x$. As observed there, the distance from $\Pi^{\prime}$ to $x$ is at most $A$, defined in equation (3.1.1) of [14] by $\cosh A=\sqrt{\frac{2}{3}(\cosh\ell_{1}+1)}$ (the same “$A$” from Lemma 4.1 here). This thus bounds the distance from $\Pi^{\prime}$ to $\tilde{\lambda}_{1}$ above.
@@ 6.1. Topological background
We now review some definitions from [8]. If $S$ is a closed surface in the interior of a $3$-manifold $M$, we denote by $M\setminus\setminus S$ the manifold with boundary obtained by splitting $M$ along $S$: it is the completion of the path metric on $Z\doteq M-S$ induced by the restriction of a metric on $M$. The inclusion map $Z\hookrightarrow M$ extends to a map $M\,\backslash\backslash\,S\to M$ that restricts on $(M\,\backslash\backslash\,S)-Z$ to a two-sheeted covering map to $S$. The restriction is a disconnected cover if and only if $S$ is two-sided in $M$.
@@ 6.2. Background on least-area surfaces
If $S$ is an oriented smooth manifold and $M$ is a Riemannian manifold, possibly with boundary, with $\dim M\geq\dim S$, any smooth immersion $f\colon\thinspace S\to M$ pulls the Riemannian metric $g$ on $M$ back to a Riemannian metric $f^{*}g$ on $S$. For any (measurable) $A\subset S$, we define the volume of $A$ under $f$, denoted by $\mathop{\rm vol}_{f}(A)$, to be the integral of the volume form of $f^{*}g$ over $A$. This extends naturally to a definition of $\mathop{\rm vol}_{f}(A)$ for any smooth map $f:S\to M$, by taking the volume form to be 00 at points where the derivative of $f$ is singular. Thus if $f$ is not an immersion, we may have $\mathop{\rm vol}_{f}(A)=0$ even if (say) $A$ has non-empty interior.
@@
Now suppose that $\|d\kappa_{\mathbf{x}}\|\leq 1$ for every $\mathbf{x}\in M_{1}$. Then for any oriented smooth manifold $S$, any smooth map $f:S\to M_{1}$, and any point $\mathbf{u}\in S$ such that $df_{\mathbf{u}}$ is non-singular, the norms defined on $T_{\mathbf{u}}$ by the pullbacks via $f$ and $\kappa\circ f$ of the metrics $g_{1}$ and $g_{2}$ on $M_{1}$ and $M_{2}$ satisfy $\|(\kappa\circ f)^{*}(g_{2})\|\leq\|f^{*}(g_{2})\|$. This implies in particular that if $\beta_{1}$ and $\beta_{2}$ denote the volume forms of the respective pullbacks, and if for $i=1,2$ we choose an $n$-form $\omega$ defining the orientation of $S$ and write $\beta_{i}=h_{i}\omega$ for some function $h_{i}$ which is positive wherever $df_{\mathbf{u}}$ is non-singular, then $h_{2}\leq h_{1}$. If as in 6.2.1 we extend the $\beta_{i}$ to all of $S$ by defining them to be 00 at points where $df_{\mathbf{u}}$ is singular, and if we define the $h_{i}$ to be 00 at such points as well, then the inequality $h_{2}\leq h_{1}$ holds on all of $S$; upon integrating we conclude that for any (measurable) set $A\subset S$ we have $\mathop{\rm vol}_{\kappa\circ f}(A)\leq\mathop{\rm vol}_{f}(A)$, in the notation of 6.2.1.
@@
Let $i:S\to M$ denote the inclusion map, and let $p\colon\thinspace\widetilde{M}\to M$ denote the covering space determined by $i_{\sharp}(\pi_{1}(S))$. Then $i$ admits a lift $\widetilde{i}:S\to\widetilde{M}$; set $\widetilde{S}=\widetilde{i}(S)$. If we write $\widetilde{M}=H^{3}/\Gamma_{0}$, where $\Gamma_{0}$ is a discrete, torsion-free group of isometries of ${\mathbb{H}}^{3}$, and let $q:{\mathbb{H}}^{3}\to\widetilde{M}$ denote the quotient map, then $\Gamma_{0}$ leaves the plane $\Pi\doteq q^{-1}(\widetilde{S})$ invariant, preserves some component of $p^{-1}(S)$, which is a plane $\Pi$, and therefore commutes with the orthogonal projection from ${\mathbb{H}}^{3}$ to $\Pi$. Hence this projection induces a projection $\pi\colon\thinspace\widetilde{M}\to\widetilde{S}$. It follows from Lemma 6.2.3 that $\|d\pi_{\mathbf{x}}\|\leq 1$ for every $\mathbf{x}\in\widetilde{M}$.
@@
[…]
@@
[…]
@@
To prove the claim we note that since $\partial N$ is totally geodesic, the manifold $N$ is acylindrical. Hence, as pointed out in Subsection 6.1, we have $\mathop{\rm kish}N=N$; we therefore have $\mathop{\rm kish}^{0}N=N$. Theorem 4.2 of [30], applied to $N$, asserts for the rightmost inequality above that “equality holds only if $N$ is decomposed into $T^{n}(0)$’s”, where in this case (for $n=3$), the truncated regular simplex $T^{3}(0)$ of edgelength 00 is a regular ideal octahedron (compare [30, Example 5.1]). But any complete manifold that decomposes into copies of $T^{3}(0)$ has cusps, and the present $N$ is compact by hypothesis. Therefore the volume inequality is strict as claimed. ∎
@@
Let $S$ be a connected, incompressible (closed) surface in (the interior of) a compact, orientable hyperbolic $3$-manifold $N$ with connected totally geodesic boundary. Then one of the following alternatives holds:
@@
[…]
@@
[…]
@@
To prove 6.2.7.3, note that since $N$ is boundary-irreducible, the surface $\partial N_{0}$ is incompressible; a priori this implies that, using a base point in $\partial N_{0}$ lying over the chosen base point of $\partial N$, the group $\pi_{1}(\widetilde{DN})$ is canonically identified with a free product with amalgamation $\pi_{1}(N_{0})\star_{\pi_{1}(\partial N_{0})}\pi_{1}(Z)$. In particular, the inclusion homomorphisms from $\pi_{1}(N_{0})$, $\pi_{1}(Z)$ and $\pi_{1}(\partial N_{0})$ to $\pi_{1}(\widetilde{DN})$ are injective, and if $A$, $B$ and $C$ denote the respective images of these injections we have $A\cap B=C$. But by the construction of $\widetilde{DN}$ we have $A=\pi_{1}(\widetilde{DN})$, and hence $B=C$; that is, the inclusion homomorphism $\pi_{1}(\partial N_{0})\to\pi_{1}(Z)$ is an isomorphism. But $\widetilde{DN}$ is aspherical since $DN$ is, and since $\partial N_{0}$ is incompressible, $Z$ is also aspherical. The genus-$2$ surface $\partial N_{0}$ is also aspherical. This implies 6.2.7.3.

### v2

@@ 3.2. Collars
We now turn our attention to embedding collars, and controlling their interaction with muffins. The definition of collar given below is implicitly used in [27] and [14]. To motivate it, we recall from Section 2.1 that for a hyperbolic $3$-manifold NN with totally geodesic boundary, we take the universal cover $\widetilde{N}$ of $N$ to be a convex subset of $\mathbb{H}^{3}$ bounded by a collection of geodesic hyperplanes. For a component $\Pi$ of $\partial\widetilde{N}$, let $\pi$ be the orthogonal projection to $\Pi$ from Definitions and Remarks 3.1.2. Then for $x\in\Pi$, $\pi^{-1}(x)$ is a geodesic intersecting $\Pi$ orthogonally at $x$, and $\pi^{-1}(x)\cap\widetilde{N}$ is either a ray with endpoint $x$ or a segment with one endpoint at $x$ and the other on a different component of $\partial\widetilde{N}$.
@@
Let $N$ be a hyperbolic $3$-manifold with totally geodesic boundary. If $x$ is a point of $\partial N$ such that there exists a (necessarily unique) geodesic path in $N$ (parametrized by arclength) which begins at $x$, is perpendicular to $\partial N$ at its initial point, and has terminal point in $\partial N$, we shall denote the length of this path by $h_{x}$, the interval $0,h_{x}]$ by $J_{x}$, and the path itself by $\alpha_{x}:J_{x}\to N$. If $x$ is a point of $\partial N$ for which no path of this type exists, then there is a unique geodesic ray beginning at $x$; in this case, we shall set $h_{x}=+\infty$ and $J_{x}=[0,\infty)$, and denote (the arclength parametrization of) the ray by $\alpha_{x}:[0,\infty)\to N$. For a subset $S$ of $\partial N$, we define the open collar of $S$ in $N$ with height $h>0$ to be the set $\bigcup_{x\in S}\alpha_{x}([0,\min(h,h_{x})))$. We shall say that the collar is embedded if $h\leq\inf\{h_{x}\,|\,x\in S\}$, and the map from $S\times[0,h]$ to $N$ defined by $(x,t)\mapsto\alpha_{x}(t)$ is one-to-one.
@@
The claim implies that the height-$h$ open collar of the preimage in $\Pi$ of $\partial N-(C\cup C^{\prime})$ is contained in $\widetilde{N}$ and does not intersect the height-$h$ open collar of any other component of $\partial\widetilde{N}$. Therefore the universal cover induces an embedding of the height-$h$ open collar of $\partial N-(C\cup C^{\prime})$ in $N$.
We now recall from Fact 3.1.7 that $U_{1}=A+W$. Here $U_{1}$, defined in Lemma 3.1.5, is a lower bound on the distance between a lift of $\lambda_{1}$ and any component of $\partial\widetilde{N}$ not containing either of its endpoints. Since $\mathrm{Muf}_{\ell_{1}}$ is contained in the $W$-neighborhood of its center, and the collar height $h$ is less than $A$, the height-$h$ open collar of $\Pi$ in $\widetilde{N}$ thus does not intersect any lift of $M$ that has no cap on $\Pi$. This implies that the height-$h$ open collar of $\partial N-(C\cup C^{\prime})$ in $N$ is disjoint from $M$. ∎
@@
Recall from the proof of Lemma 3.1.8 that $\widetilde{M}$ is contained in a radius-$W_{2}$ neighborhood of its center $\tilde{\lambda}_{2}$, where $W_{2}$ is its waist radius, and from Lemma 3.1.5 that the quantity $U_{2}$ defined in that lemma is a lower bound for the distance from $\tilde{\lambda}_{2}$ to $\Pi_{3}$. From the proof of Proposition 3.1.11, for $(x,y)$ in the rectangle $a,b]\times[c,d]$, $W_{2}$ is bounded above by $W^{0}_{S}$ and $U_{2}$ is bounded below by $U_{2}^{0}$. Therefore by the triangle inequality, no point lies in both $\widetilde{M}$ and a height-$H^{0}$ open collar of $\Pi_{3}$. ∎
@@ 4. Volume bounds
For the muffin $\mathrm{Muf}_{\ell_{1}}$ from Definitions and Remarks 3.1.6, depending on a parameter $\ell_{1}$, let $\mathit{VM}(x)$ record its volume as a function of $x=\cosh\ell_{1}$. Then
(4.1.1) $\mathit{VM}^{\prime}(x)=2\pi A\frac{d}{dx}(\cosh R)<0$
@@
$\mathit{VM}^{\prime}(x)=\pi\left[(2A)[\cosh R]^{\prime}+(2A)^{\prime}\cosh R-\frac{1}{\sqrt{x^{2}-1}}\right]=2\pi A[\cosh R]^{\prime}$
@@
(4.2.1) $\mathrm{vol}(N)\geq\mathit{VM}(x)+\pi(2-\cosh R)(2H+\sinh(2H)),$
where $\mathit{VM}(x)$ records the volume of $\mathrm{Muf}_{\ell_{1}}$ as in Lemma 4.1, and for $A$ equal to the altitude length of $\mathrm{Muf}_{\ell_{1}}$ as described there, $H=\min\{A,\ell_{2}/2\}$.
@@
$\frac{\partial V}{\partial x}(x,H)=\mathit{VM}^{\prime}(x)-\pi[\cosh R]^{\prime}(2H+\sinh(2H))=\pi[\cosh R]^{\prime}\left(2A-2H-\sinh(2H)\right).$
Since $\cosh R$ decreases with $x$, this quantity is positive as long as $2A<2H+\sinh(2H)$. For $x\leq 1.5$ we have $2A<\cosh^{-1}(2.34)<1.5$ and for $H\geq\cosh^{-1}(1.5)/2$,
$2H+\sinh(2H)\geq\cosh^{-1}(1.5)+\sqrt{1.5^{2}-1}>2.08.$
@@
Table 1. Values of $V(x,H)$, truncated after 33 decimal places, for $H=\cosh^{-1}(Y(x))/2$.
@@
$a\leq\cosh\ell_{1}\leq b\leq 2\quad\mbox{and}\quad c\leq\cosh\ell_{2}\leq d\leq 3.$
For $W_{R}^{0}$, $W^{0}_{S}$, $T_{12}^{0}$ and $T_{22}^{0}$ as in (3.1.11), if $W^{0}_{R}+W^{0}_{S}\leq T^{0}_{12}$ and $2W_{S}^{0}\leq T^{0}_{22}$, then $\mathrm{vol}(N)\geq\mathit{VM}(b)+\mathit{VM}_{2}^{0}$, where $\mathit{VM}(x)$ is as in Lemma 4.1 and $\mathit{VM}_{2}^{0}$ is defined in (4.3.1).
@@
(4.4.1) $\mathrm{vol}(N)\geq\mathit{VM}(b)+\mathit{VM}_{2}^{0}+\pi(3-\cosh R(a)-\cosh S(b,c))\left(2H+\sinh(2H)\right),$
@@
Recall from Lemma 4.1 that the function $\mathit{VM}(x)$, which records the volume of $\mathrm{Muf}_{\ell_{1}}$ as a function of $x=\cosh\ell_{1}$, is decreasing. Thus since $x\leq b$, $\mathit{VM}(x)\geq\mathit{VM}(b)$. By Lemma 4.3, $\mathit{VM}_{2}^{0}$ bounds the volume of $M(\ell_{2},S)$ below for these values of $\ell_{1}$ and $\ell_{2}$. Thus $\mathrm{vol}(N)\geq\mathit{VM}(b)+\mathit{VM}_{2}^{0}$.
@@
On each maximal subinterval $a_{0},b_{0})$ of $[1.24,1.5)$ where $Y$ is constant, we find the minimum value of the function $\max\{\ell_{1},E,M\}$ from Lemma 2.2.2. Let $c_{0}$ be the hyperbolic cosine of this value, truncated after three decimal places. By Lemma 2.2.2, for $N$ with $\cosh\ell_{1}\in[a_{0},b_{0}]$, $\cosh\ell_{2}\geq c_{0}$. Let $d_{0}$ be the value of $Y$ on $a_{0},b_{0})$. We have written Python scripts that divide the rectangle $a_{0},b_{0}]\times[c_{0},d_{0}]$ into subrectangles of size $0.001\times 0.001$, and on each subrectangle $a,b]\times[c,d]$ performs the following tasks.
@@
    Check whether $a,b]\times[c,d]$ intersects the possible subregion: the subset of $a_{0},b_{0}]\times[c_{0},d_{0}]$ consisting of $(x,y)$-values that satisfy the criterion of Lemma 2.2.2. For instance, $E(x)=\max\{\ell_{1},E,M\}$ for $x\leq 1.366$, so since $E$ is a decreasing function of $x$ here (cf. [14, Lemma 3.4]), for $b\leq 1.366$ this is tantamount to checking that $d\geq\cosh E(b)$.
    If $a,b]\times[c,d]$ does not intersect the possible subregion then we ignore it since in this case, by Lemma 2.2.2, there does not exist a manifold satisfying the hypotheses of the present result and with an $(x,y)$-value in $a,b]\times[c,d]$.
@@
    If $a,b]\times[c,d]$ does intersect the possible subregion, check the muffin-embedding criteria of Proposition 3.1.11. These hold in each such case, confirming the expectation recorded above that result.
@@
    On $a,b]\times[c,d]$ intersecting the possible subregion, compute the two volume bounds supplied by Proposition 4.4, recording the larger. (The reason that the sum of muffin volumes $\mathit{VM}(b)+\mathit{VM}^{0}_{2}$ may be larger than the bound (4.4.1), which also incorporates the volume of a collar of a region in $\partial N$, is that in some cases the given lower bound $3-\cosh R(a)-\cosh S(b,c)$ for the area of that region may be negative.)
@@
The resulting volume bounds are summarized in Figure 4.2. Each big rectangle $a_{0},b_{0}]\times[c_{0},d_{0}]$ from above is shaded in the Figure, and the smallest of the volume bounds produced by the process above—taken over all subrectangles $a,b]\times[c,d]$—is recorded in or directly above the big rectangle. (The subrectangle on which the minimum is attained is also blackened in the Figure.) Since each of these bounds is larger than $7.4$, the Theorem follows from Proposition 4.2. ∎
@@
We use a single large rectangle $1.23,1.24]\times[1.74,1.938]$. The lower bound of $y=1.74$ is the minimum of $\cosh E(x)$ on this interval, which is attained at $x=1.24$, truncated after three decimal places. The upper bound is the maximum of $Y_{0}(x)$ on the interval, also attained at $1.24$, rounded up after three decimal places. Because $Y_{0}(x)$ is not constant here but increasing, we alter step (1) above for a subrectangle $a,b]\times[c,d]$ by checking that both $d\geq\cosh E(b)$ and $c\leq Y_{0}(x)$.
To achieve a slight improvement in our volume bounds, we use a subrectangle size of $0.0005\times 0.001$—that is, we divide the previous subrectangles in half vertically. The smallest bound resulting from Theorem 4.5’s revised process is $7.409$, truncated after three decimal places, attained on the subrectangle $1.2305,1.231]\times[1.808,1.809]$, proving the result. (Using the previous subrectangle size would yield $7.399$ as a bound.) The relevant Python script is VolScript_EUl2.py, and its output is 123_124.txt. ∎
@@ 5. Trimonic Manifolds from a $(1,1,1)$-hexagon
As in the previous proof, we take $\Pi^{\prime}$ to be the component of $\partial\widetilde{N}$ containing the external edge of $C$ closest to $x$. As observed there, the distance from $\Pi^{\prime}$ to $x$ is at most $A$, defined in equation (3.1.1) of [14] by $\cosh A=\sqrt{\frac{2}{3}(\cosh\ell_{1}+1)}$ (the same “AA” from Lemma 4.1 here). This thus bounds the distance from $\Pi^{\prime}$ to $\tilde{\lambda}_{1}$ above.
@@ 6.1. Topological background
[…]
@@ 6.2. Background on least-area surfaces
If $S$ is an oriented smooth manifold and $M$ is a Riemannian manifold, possibly with boundary, with $\dim M\geq\dim S$, any smooth immersion $f\colon\thinspace S\to M$ pulls the Riemannian metric $g$ on $M$ back to a Riemannian metric $f^{*}g$ on $S$. For any (measurable) $A\subset S$, we define the volume of $A$ under $f$, denoted by $\mathop{\rm vol}_{f}(A)$, to be the integral of the volume form of $f^{*}g$ over $A$. This extends naturally to a definition of $\mathop{\rm vol}_{f}(A)$ for any smooth map $f:S\to M$, by taking the volume form to be 0 at points where the derivative of $f$ is singular. Thus if $f$ is not an immersion, we may have $\mathop{\rm vol}_{f}(A)=0$ even if (say) $A$ has non-empty interior.
@@
Now suppose that $\|d\kappa_{\mathbf{x}}\|\leq 1$ for every $\mathbf{x}\in M_{1}$. Then for any oriented smooth manifold $S$, any smooth map $f:S\to M_{1}$, and any point $\mathbf{u}\in S$ such that $df_{\mathbf{u}}$ is non-singular, the norms defined on $T_{\mathbf{u}}$ by the pullbacks via $f$ and $\kappa\circ f$ of the metrics $g_{1}$ and $g_{2}$ on $M_{1}$ and $M_{2}$ satisfy $\|(\kappa\circ f)^{*}(g_{2})\|\leq\|f^{*}(g_{2})\|$. This implies in particular that if $\beta_{1}$ and $\beta_{2}$ denote the volume forms of the respective pullbacks, and if for $i=1,2$ we choose an $n$-form $\omega$ defining the orientation of $S$ and write $\beta_{i}=h_{i}\omega$ for some function $h_{i}$ which is positive wherever $df_{\mathbf{u}}$ is non-singular, then $h_{2}\leq h_{1}$. If as in 6.2.1 we extend the $\beta_{i}$ to all of $S$ by defining them to be 0 at points where $df_{\mathbf{u}}$ is singular, and if we define the $h_{i}$ to be 0 at such points as well, then the inequality $h_{2}\leq h_{1}$ holds on all of $S$; upon integrating we conclude that for any (measurable) set $A\subset S$ we have $\mathop{\rm vol}_{\kappa\circ f}(A)\leq\mathop{\rm vol}_{f}(A)$, in the notation of 6.2.1.
@@
[…]
@@
In the proof of Lemma 6.2.5 and below it, the double of a manifold $N$ with boundary is the manifold $\mathit{DN}$ obtained from $N\sqcup\overline{N}$, where $\overline{N}$ is a second copy of $N$, by identifying $\partial N\to\partial\overline{N}$ via the identity map. The following facts are standard and will be taken for granted. First, if $N$ is oriented then so is $\mathit{DN}$, by equipping $\overline{N}$ with the opposite orientation from that of $N$. Second, a hyperbolic structure with totally geodesic boundary on $N$ embeds isometrically into a boundaryless hyperbolic structure on $DN$ in which $\partial N$ hence sits as totally geodesic surface separating $N$ from $\overline{N}$.
@@
Theorem 7.2 of [4] asserts that $\mathop{\rm vol}N\geq\frac{1}{2}V_{3}\|\mathit{DN}\|$, where $V_{3}$ is the volume of a regular ideal tetrahedron and $\|\cdot\|$ denotes the Gromov norm.
As used in the proof of [4, Th. 9.1], $\mathop{\rm kish}^{0}N$ admits a hyperbolic structure with totally geodesic boundary, which by Miyamoto’s universal lower bound on the volumes of manifolds with totally geodesic boundary [30, Th. 4.2] has volume at least $V_{\rm oct}\,\bar{\chi}(\mathop{\rm kish}N)$. Its double $D\mathop{\rm kish}^{0}N$ therefore has a finite-volume hyperbolic structure with volume at least twice this. And $D\mathop{\rm kish}^{0}N$ is the interior of the submanifold $D\mathop{\rm kish}N$ of $\mathit{DN}$ whose frontier is a disjoint union of incompressible tori. It thus follows that $V_{3}\|\mathit{DN}\|\geq\mathop{\rm vol}(D\mathop{\rm kish}^{0}N)$ from Theorem 6.5.5 of Thurston’s notes [35]. Combining these inequalities, we obtain
$\mathop{\rm vol}N\geq\frac{1}{2}V_{3}\|\mathit{DN}\|\geq\frac{1}{2}\mathop{\rm vol}(\textstyle{D\mathop{\rm kish}^{0}}N)\geq V_{\rm oct}\,\bar{\chi}(\mathop{\rm kish}N)$
@@
To prove the claim we note that since $\partial N$ is totally geodesic, the manifold $N$ is acylindrical. Hence, as pointed out in Subsection 6.1, we have $\mathop{\rm kish}N=N$; we therefore have $\mathop{\rm kish}^{0}N=N$. Theorem 4.2 of [30], applied to $N$, asserts for the rightmost inequality above that “equality holds only if $N$ is decomposed into $T^{n}(0)$’s”, where in this case (for $n=3$), the truncated regular simplex $T^{3}(0)$ of edgelength 0 is a regular ideal octahedron (compare [30, Example 5.1]). But any complete manifold that decomposes into copies of $T^{3}(0)$ has cusps, and the present $N$ is compact by hypothesis. Therefore the volume inequality is strict as claimed. ∎
@@
Let $S$ be a connected, incompressible (closed) surface in (the interior of) a compact, orientable hyperbolic $3$-manifold NN with connected totally geodesic boundary. Then one of the following alternatives holds:
@@
Proposition 6.2.6, applied with the double $\mathit{DN}\supset N$ playing the role of $M$, implies that the inclusion $S\hookrightarrow\mathit{DN}$ is homotopic in $\mathit{DN}$ to a least-area immersion $f\colon\thinspace S\to\mathit{DN}$ which is either (i) an embedding, or (ii) a two-sheeted covering map to a one-sided surface $K\subset DN$. (The second alternative of the conclusion of Proposition 6.2.6 is stronger than (ii), but the stronger information will not be used.)
The inclusion map $\partial N\hookrightarrow\mathit{DN}$ of the totally geodesic surface $\partial N$ is least-area, by Corollary 6.2.4, so since $S$ and $\partial N$ are disjoint, Theorem 6.2 of [16] further asserts that $f(S)$ is also disjoint from or identical to $\partial N$. In view of 6.2.7.1 we cannot have $f(S)=\partial N$. Hence:
@@
Let us now fix a base point in $\partial N\subset N\subset DN$, and consider the covering space $p\colon\thinspace\widetilde{\mathit{DN}}\to\mathit{DN}$ determined by the image of the inclusion homomorphism $\pi_{1}(N)\to\pi_{1}(DN)$. There is a submanifold $N_{0}$ of $\widetilde{DN}$ which is mapped diffeomorphically onto $N$ by $p$; and since $\partial N$ is connected, the submanifold $Z\doteq\widetilde{DN}-\mathop{\rm int}N_{0}$ is also connected. We claim:
@@
[…]